[7] A3. (a) Draw the streamlines and vortex lines of a Rankine vortex. Indicate which field lines are streamlines and which f could you please help me with answering all parts of this question. like and comment are rewarded.
Answer #1

a) forced vorted 7a Re EX Potential (free) velocity distance Core radius

One of the most features of the Rankine vortex is its vorticity field. actually consistent with
the definition of the vortex velocity field (1.1) and therefore the the appliance of the curl operator
in cylindrical coordinates (A 4), it's evident that the vortex presents vertical vorticity
component only. Furthermore the vorticity field modulus is constant within the inner a part of
the vortex, it's positive and it's a function of the utmost flow velocity and therefore the vortex
characteristic distance only. within the outer region of the vortex, the flow has no vorticity
at all.

V xv = k 1 arve) kl ve ave + 24. if 0<r<R 0 if R<r где ar

It is worth to notice that the Rankine vortex is characterized by endless velocity
field, but with a discontinuity in vorticity at the characteristic distance.

b) Let u(t,x) represent the speed vector field of the fluid. Let x(t) denote the position of a particle moving with the fluid, then the speed x˙(t) of the particle at a time t are going to be adequate to the speed of the fluid flow at the purpose (t,x(t)), namely
u(t,x(t))=x˙(t)
Now suppose that u(t,x)⋅∇H(t,x)=0. He want to point out that this suggests that H is constant along the trail of a particle moving with he fluid. Notice that for any path x(t) we've
d Ht, x(t) dt = aH (t, x(t)) + X(t) VH(t, x(t)) at
Assuming then that ∂tH=0, and assuming that the trail x(t) is that of a particle moving with he fluid, the equations written above imply
d - H (t, x(t)) = u(t, x(t)) · JH (t, x(t)) = 0 dt
so the quantity H is constant along a flow line, as desired!

c)Re: fv.d хKхбочx tox (1-2-) xo2. 1-охов =k P= Constant density of viscous fund 2-xxlo 3 x 2 16о о Re 160xE у : х. = (, ).л